A body is projected downward at an angle of 20° with respect to the horizontal(+X axis), from the top of a building 170 m high. Its initial speed is 40 m/s. (a) How long will it take before striking the ground?


1
Expert's answer
2022-02-06T22:38:01-0500

The initial vertical velocity "= u \times sin 20^0 = 40\times sin 20^0 = 13.68 m/s."

The vertical height given (h) = 170 meters

Using the formula "h = ut + \frac{1}{2} gt^2"

"170 = 40\times t + 4.9t^2"

Solving the quadratic equation, we get

t= 3.08 seconds.

The range is given by "x= u cos\theta * t"

"x= 40 * cos 20^0 * 3.08"

Solving,

"x=115.77" meters.